1, Suppose the baseline risk for a disease is 22%. However, researchers find that regularly riding on the bus increases risk by 31%.

A) For those who regularly ride the bus, how many total people out of 1000 are at risk for this disease?

Ans: A total of  out of 1000 are at risk.

B) How many additional people above baseline are at risk?

Ans: An additional  people out of 1000 above the baseline are at risk.

2, The U.S., with a population of about 300 million, consumes about 20 million barrels of oil a day. India has a population of 1.2 billion people. If they consumed oil at the same rate or proportion as the U.S. how much oil would they consume?

    million barrels of oil a day
Answer #1

A.

Those who ride bus regularily, 31% are at risk

which means 31 out of 100 people are at risk,

Out of 1000 people 310 people are at risk

B.

Base line risk% is 22%

Out of 1000 people people at baseline risk = 22/100*1000 = 220

Additional people = 310 - 220 = 90

2.

Indias population is 1200 million, and US population is 300 million

Indias population is 1200/300 = 4 times US population

US consumes 20 million barrels of oil

India will consume = 20*4 = 80 million barrels of oil